10) How many grams of Pbar would be deposited by the electrolysis of a Pbsoa solution at 32 amps for 2 hours?
Answer #1

Number of electrons = I*t / F

= 3.2 amp * 2 * 3600 s / (96485)

=0.2387 mol electrons

2 moles of electrons are required for 1 mole of Pb

So moles of Pb= 0.5 *0.2387 = 0.1193 moles

Mass of Pb= 0 .1193* 207.2 = 24.74 g of Pb